Lucido Products markets two computer games: Claimjumper and Makeover. A contribution format income statement for a recent month for the two games appears below:

Claimjumper Makeover Total
Sales $ 116,000 $ 58,000 $ 174,000
Variable expenses 28,840 5,960 34,800
Contribution margin $ 87,160 $ 52,040 139,200
Fixed expenses 88,560
Net operating income $ 50,640

Required:

1. What is the overall contribution margin (CM) ratio for the company?

2. What is the company's overall break-even point in dollar sales?

3. Prepare a contribution format income statement at the company's break-even point that shows the appropriate levels of sales for the two products.
Answer #1

1. Contribution margin (CM) ratio = Contribution /Sales = $139,200 / $174,000 = 80%

2. Break-even point in dollar sales = Fixed expenses / Contribution margin ratio
= $88,560 / 0.80 = $110,700

3.

Claimjumper Makeover Total
Sales 116,000/174,000 x 110,700 = $73,800 58,000/174,000 x 110,700 = $36,900 $110,700
Variable expenses 28,840/34,800 x 22,140= $18,348 5,960/34,800 x 22,140= $3,792 $110,700 x .20 = $22,140
Contribution margin $55,452 $33,108 $88,560
Fixed expenses $88,560
Net operating income $0